The perimeter of the yard is 130 feet. If the length is 10 less than 2 times the width, what are the dimensions? Show all work.
Thanks alot !

Answer by checkley75(3666)

L=2W-10 IS THE LENGTH
W=WIDTH
2W+2(2W-10)=130
2W+4W-20=130
6W=130+20
6W=150
W=150/6
W=25 IS THE WIDTH.
2*25-10=50-10=40 IS THE LENGTH.
PROOF
2*25+2*40=130
50+80=130
130=130